OpenStudy (anonymous):

what is the different of representation of matrix of a)T(1,x,x^2)={1+x ,3-x^2,4+2x-3x^2} b)(x,y,z)=(x+y+z,x-y-3z,-x+y+3z) c){(0,0,1),(1,0,1),(0,0,1)}

OpenStudy (anonymous):

@TuringTest help pls

OpenStudy (turingtest):

What is the question asking for exactly?

OpenStudy (anonymous):

how to represent this matrix are the the same or different?

OpenStudy (turingtest):

well starting with the first one, we should write out what we know\[A\vec x=\vec b\]\[A\left[\begin{matrix}1\\x\\x^2\end{matrix}\right]=\left[\begin{matrix}1+x\\3-x^2\\4+2x-3x^2\end{matrix}\right]\]

OpenStudy (turingtest):

this tells you that \(A\) is going to be a \(3\times3\) matrix because the original and final vectors are \(3\times1\), now you just need to identify the elements that the matrix needs to have

OpenStudy (turingtest):

what must the first row of A be?

OpenStudy (turingtest):

\[A_1\vec x=\vec b\]\[\left[\begin{matrix}a_{11}&a_{12}&a_{13}\end{matrix}\right]\left[\begin{matrix}1\\x\\x^2\end{matrix}\right]=\left[\begin{matrix}1+x\end{matrix}\right]\]

OpenStudy (turingtest):

^that is the first row of A did you try refreshing yet?

OpenStudy (anonymous):

we will have this for number a